Sarah and James Hernandez purchased 210 shares of Macy’s stock at $38 a share. One year later, they sold the stock for $70.00 a share. They paid a broker a commision of $7 when they purchased the stock and a commision of $11 when they sold the stock. During the 12-month period the couple owned the stock, Macy’s paid dividends that totaled $1.72 a share. Calculate the Hernandez total return for this investment.
Answer #1


Solution: Sale value=210*70 Add : Dividend income=210*1.72 Less : Buy value=-210*38 Less : commission paid=-(7+11) total retu